1 point) How much would need to be invested today in order to have $14500 in 18 months time if interest can be earned at 5% compounded continuously? Answer: $
Answer #1

Interest is compounded continuosly means each period interest is added to principal and interest is not only on principal but on interest portion also.

Present Value with Continuous Compounding = Future Value / e^(rate*time)

= 14500/e^(.05/12 * 18)

= 14500/e^0.075

= 14500/1.07788415088

= 13452.2805518

= $13,452.28